Large Cardinal Results and Problems

Richard Laver

We will survey some results on very large cardinals, including facts about the axioms themselves and about some algebraic and other low-level statements which arise in the study of this topic.

Address: Department of Mathematics, University of Colorado, Boulder, Colorado, 80309, USA

E-mail: laver@euclid.colorado.edu